A fair spinner has 9 equal sections 3 red 2 blue 4 green it is span twice what’s the probability of getting blue both times.


1
Expert's answer
2021-02-24T12:20:24-0500

Given "P(red)=\dfrac{3}{9}, P(blue)=\dfrac{2}{9}, P(green)=\dfrac{4}{9}"

Then


"P(blue\ and\ blue)=\dfrac{2}{9}(\dfrac{2}{9})=\dfrac{4}{81}"

The probability of getting blue both times is "\dfrac{4}{81}."